Please help me find the equation of this function problem: Assume that it costs $0.30 to mail a letter weighing one ounce or less, and then $0.27 for each additional ounce or fraction of an ounce. Let L(x)be the cost of mailing a letter weighing x ounces. Find L(4).
Found 2 solutions by edjones, oberobic:
Answer by edjones(7311)

L(4)=0.3+3*.27
=.3+.81= $1.11
Ed

Answer by oberobic(2302)

Think of this as a linear equation.
y = mx + b is the general equation for a line.
We're told that for each ounce after the first ounce, the charge is 27 cents. That is the slope (m).
We're told the first oz costs 30 cents = 27+3.
y = 27x + 3
Checking at x = 1,
y = 27+3 = 30.
Checking at x = 2,
y = 27*2 + 3 = 54 + 3 = 57.
That fits the facts given.
So, apply for 4 oz:
y = 27*4 + 3 = 108+3 = 111 cents = $1.11.